# Securing WiFi Fingerprint-based Indoor Localization Systems from Malicious Access Points

### v1

@@ Abstract
WiFi fingerprint-based indoor localization schemes deliver highly accurate location data by matching the received signal strength indicator (RSSI) with an offline database using machine learning (ML) or deep learning (DL) models. However, over time, RSSI values degrade due to the malicious behavior of access points (APs), causing low positional accuracy due to RSSI value mismatch with the offline database. Existing literature lacks detection of malicious APs in the online phase and mitigating their effects. This research addresses these limitations and proposes a long-term reliable indoor localization scheme by incorporating malicious AP detection and their effect mitigation techniques. The proposed scheme uses a Light Gradient-Boosting Machine (LGBM) classifier to estimate locations and integrates simple yet efficient techniques to detect malicious APs based on online query data. Subsequently, a mitigation technique is incorporated that updates the offline database and online queries by imputing stable values for malicious APs using LGBM Regressors. Additionally, we introduce a noise addition mechanism in the offline database to capture the dynamic environmental effects. Extensive experimental evaluation shows that the proposed scheme attains a detection accuracy above 95% for each attack type. The mitigation strategy effectively restores the system’s performance nearly to its original state when no malicious AP is present. The noise addition module reduces localization errors by nearly 16%. Furthermore, the proposed solution is lightweight, reducing the execution time by approximately 94% compared to the existing methods.
Index Terms:
WiFi fingerprint, indoor localization, security, malicious API Introduction
With the rapid advancement of wireless communication technologies and the Internet of Things (IoT), location-based services have become indispensable in many application domains. Various localization mechanisms that operate either indoors or outdoors are employed to determine location. In outdoor environments, positions can be accurately determined using the Global Navigation Satellite System (GNSS). However, indoor positioning is challenging due to the attenuation of the Global Positioning System (GPS) signal caused by building structures and their materials [1]. Accurate indoor location tracking is critical for several purposes such as firefighter tracking under hazardous conditions and in emergencies, location-based smart resource management, robotics, shopping mall navigation, healthcare, and virtual reality [1], [2], [3]. Consequently, indoor localization has become a major focus for both academia and industry.
@@
A significant security issue is the presence of malicious access points that generate false RSSI signals [8]. Over time, APs may behave maliciously due to component malfunctions [9, 10] and security attacks such as signal jamming [8], spoofing [6], distance fraud [11], [12], and physical attacks [5]. These malicious APs mislead localization systems and reduce location accuracy. Unreliable location accuracy can have serious consequences in critical situations. For example, if a person needs immediate medical attention in a hospital, inaccurate location prediction of healthcare professionals and facilities can elevate health risks. Similarly, robots navigating in complex indoor environments with unreliable localization methods may cause damage to assets and potentially lead to injuries. Existing literature deals with malicious APs during the initial database construction phase (offline phase) by selecting a subset of feasible APs [5], [6]. However, filtering out malicious APs only in the offline phase does not guarantee the long-term robustness of the indoor localization systems as a reliable AP may become malicious over time. Current literature lacks research on detecting the long-term malicious behavior of APs [10, 13, 14, 15].
After detecting malicious APs, it is crucial to mitigate their effect to ensure the accurate functioning of the indoor localization system. One approach to achieving this goal is to reconstruct the database. However, each time a malicious AP is detected, conducting a site survey to build the database in a dynamic environment is highly expensive and labor-intensive [13]. Existing literature has explored the use of crowdsourcing techniques [15] for creating and updating the fingerprint database, which significantly reduces human labor. However, this scheme is vulnerable to various security issues [16]. Additionally, robots are utilized for autonomous database construction, which can be expensive and may not perform effectively in a challenging environments [17]. Instead, it would be more beneficial to develop a mechanism that enables an indoor localization system to function accurately despite the presence of a reasonable percentage of malicious APs, eliminating the need for frequent database reconstruction. This would enhance the system’s accuracy while reducing the costs associated with repeated database creation. Although several existing research works mitigate the effect of RSSI value alteration due to environmental factors [18], [15], [13], none of them consider the malicious behavior of APs causing abrupt and dynamic changes to RSSI values.
@@ II Related works
WiFi RSSI-based indoor localization schemes provide highly accurate location data at low additional cost. However, the location predicted by these models often becomes erroneous due to the alteration of RSSI values caused by malicious AP behaviour [5], [11], [8], changes in layout [15], removal of AP [14], position change of AP [14], [15], and environmental factors [4, 18, 14]. Existing literary works [5, 6, 19, 20] detect and filter malicious APs during the model training phase (offline phase) of indoor localization systems. Here, a subset of optimal AP is selected, and using these selected nonmalicious APs, a fingerprint database is constructed to train the machine learning model. The selection of correlated APs increases localization accuracy and reduces computational complexity. Wang et al. [5] utilized interclass dispersion to calculate AP confidence and selected APs with high confidence. Ye et al. [6] used the Pearson Correlation Coefficient (PCC), where a lower correlation value indicates greater susceptibility of an AP to being malicious. Chen et al. [19] considered the resolution capability to select a set of APs, where APs are chosen based on information gain. Panja et al. [20] used the Binary Particle Swarm Optimization (BPSO) to select a set of suitable APs. In contrast, our proposed scheme uses Spearman’s Correlation Coefficient (SRCC) to select correlated APs in the offline phase. We use normalization, correlated AP selection, and noise addition as data preprocessing, where normalization speeds up the computational time, and noise addition makes the system robust to long-term signal variation due to dynamic environmental change.
@@
Detecting malicious APs only in the offline phase is not sufficient for long-term model reliability. Several works explore long-term model robustness challenges in the online phase. For example, Yan et al. [18] proposed a crowdsourcing-based mechanism to handle RSSI signal value alteration in the online phase. This scheme utilizes a denoising autoencoder to address RSSI alteration caused by environmental dynamics. It requires extensive high-quality crowdsourced data for training, and the model is updated frequently, which can be resource-intensive. Li et al. [10] addressed signal variation caused by dynamic environmental change and temporal effects using disagreement-based semi-supervised learning. This scheme struggles to handle sudden and significant alterations in the environment, which can impact localization accuracy. Tiku and Pasricha [4] utilized Siamese neural encoders to address RSSI signal variation without the need for model retraining. This scheme handles signal alteration caused by environmental changes, human movement, and removal or replacement of AP. Significant changes in the environment can affect localization accuracy. Huang et al. [13] addressed location accuracy degradation caused by environmental changes. This scheme utilizes a Gaussian Process Regression to create an offline database and incorporates the Marginalized Particle Extended Gaussian process that uses crowdsourced data to recursively update the database. However, this scheme is susceptible to localization errors due to rapid fluctuations in signal strength, and its effectiveness depends on the quality of the crowdsourced data. Jiang et al. [14] addressed RSSI signal alteration caused by the change in the number of AP and environmental factors. The authors proposed a feature adaptive extreme learning machine-based approach that can adapt to a changing number of features. This scheme faces difficulties to cope with sudden environmental changes and runs the risk of overfitting to recent signal variations. Yang et al. [15] proposed the Altered AP Identification and Fingerprint Updating (AAIFU) scheme, which detects altered APs in the online phase by considering changes in RSSI values due to positional changes of APs. This scheme uses a Gradient Boosting Decision Tree (GBDT) regression model to identify the positional changes of APs. It then updates the database by replacing the outdated RSSI values with fresh RSSI values obtained through GBDT regression. However, the effectiveness of this scheme depends on a large volume of crowdsourced data. Moreover, positional changes of an AP only create persistent changes in RSSI values. Among these works, Li et al. [10], Huang et al. [13], Jiang et al. [14], and Yang et al. [15] update database to keep the fingerprints fresh. On the other hand, none of the schemes except Yang et al. [15], identify the causes of RSSI signal alteration. In contrast to the existing works, our proposed scheme addresses the RSSI value alteration caused by malicious AP behavior, characterized by abrupt and random changes in RSSI values. Our proposed scheme detects malicious APs in real time by analyzing data from various location queries. Additionally, it utilizes the LGBM regression model [21] to update the existing database, which mitigates the effect of malicious APs on the performance of the indoor localization scheme. Here to note that the proposed scheme handles both static and dynamic changes in RSSI values.
@@ III System Model
As shown in Fig. 1, the proposed scheme augments the existing WiFi fingerprint-based indoor localization system [22] by including a malicious effect mitigation module and online database. The proposed system model comprises access points and reference points (RPs). The offline database comprises a set of records of form , where is the RSSI vector and is the RSSI value of th AP measured at th RP, and is the location of the th RP, , and . In the online phase, a device measures a fingerprint vector with respect to its location and sends it to the location prediction module. The location prediction module matches with the records stored in the offline database using either a ML or DL model and returns the estimated location to the mobile device. The online database is constructed by saving the results of online queries along with associated RSSI vectors and used to detect malicious APs in the online phase. The malicious effect mitigation module detects malicious APs and takes measures to diminish their effect. The proposed system model subjects to false fingerprints both in online and offline phases due to various environmental factors [4, 14], temporal effects [6], malicious behavior of APs [5, 6]. We assume that the offline and online databases, the location prediction module, and the malicious effect mitigation module reside in a trusted server.
@@ IV-A1 Database Construction
At first, a reliable database is created using site survey [23]. This mechanism involves a person with multiple reliable mobile devices visiting every RP. It collects several labeled samples in the form , where is a RSSI fingerprint vector computed at th RP and is the location of the th RP. The raw database created by site survey is a collection of records .
@@ IV-A2 Correlated AP Selection
The proposed scheme selects a set of reliable APs from the available ones in the environment using the Spearman’s Rank Correlation Coefficient (SRCC) [24] method with a threshold value set to 0.1 empirically, which is optimal for the considered UJI’s dataset [25]. Suppose the number of selected correlated APs is . The proposed scheme constructs an offline database from the raw database constructed via site survey, where contains RSSI entries only for the selected APs.
@@ IV-C1 Check Variance
It checks the variance of an AP for every RP in the online database. If the maximum variance value of an AP equals or exceeds the variance threshold, the proposed scheme marks that AP as malicious. The variance threshold is set to 0.05, as nearly 99% of the variance for each AP in the offline database is within 0.05 for the UJI dataset [25]. Let,
@@ IV-D2 Update Online Queries and Offline Database
The LGBMRegressors predicts RSSI values when input data is of form and use these data to replace the values of malicious APs , in the offline record. Similarly, when the input data is of form , the predicted RSSI values replace , in the online query . The localization model is retrained with the updated offline database. The updated query is used by the classification model to predict the location. The process of predicting RSSI values is inspired by Montoliu et al. [27].
@@ V Experimental Evaluation
We implemented the proposed scheme using Python and evaluated its performance using the UJI’s dataset [25]. The UJI’s dataset [25] is the longest public dataset with 103584 WiFi fingerprints. We used only the data of floor 3 for brevity, as also done in [4]. This data set contains 4320 data samples for training purposes for floor 3 [28]. We incorporated noise-added samples in this set and extended its size to 4800. Initially, there were 620 APs in the environment. However, we took only 40 APs after correlated AP selection. We trained our localization model which is a LGBMClassifier [21], with the offline dataset. We used seven performance metrics shown in Table III to evaluate the performance. The parameters used in defining performance metrics are shown in Table II.
@@ V-A Impact of Noise Addition
Figure 3 shows mean localization errors over the data of 25 months from UJI’s dataset [25]. The curve labeled SRCC (0.1) shows the effect of correlated AP selection only, the curve labeled SRCC (0.1) + Normalized shows the impact of AP selection and normalization, and the curve labeled SRCC (0.1) + Normalized + Noise (0.5) shows the effect of AP selection, normalization, and noise addition. Figure 3 shows a notable reduction in localization error for the curve SRCC (0.1) + Normalized + Noise (0.5) with respect to SRCC (0.1) + Normalized after the 10th month because the model becomes robust to RSSI fluctuations in the long term. The MLE is reduced by nearly 16% in the 16th month, which is the maximum change of the red curve with respect to the green one. Normalizing data helps models converge faster and adding noise makes the model more robust against long-term noise. Correlated AP selection, noise addition, and data normalization bring the MLE of the proposed scheme below 2 meters, outperforming existing machine learning schemes [4].
@@ V-B1 Performance Analysis of the Proposed Scheme
False Positive Rate (FPR): Figure 4(a) shows that the FPR consistently remains below 4%. The legitimate APs sometimes exhibit signal variation due to temporal effect and are marked as malicious by the detection system.
@@
False Negative Rate (FNR): Figure 4(b) shows that except for the ARCO attack, all other attacks show a negligible FNR. As marking an AP as malicious in our experiment was a random process, in very few cases, the selected malicious AP was assigned the same original RSSI value, leading to a false negative outcome. In ARCO attacks, FNR stays nearly 10%. In this attack, we add or subtract a fixed value and also add an offset in the noise addition module, leading to very little change in the mean that can not be distinguished with the mean difference checker, resulting in an increased FNR.
@@
Recall: The recall is consistently 100% for RV and ARRO attacks, almost 98% for CV attacks, and nearly 92% for ARCO attacks, as shown in Fig. 4(c). In ARCO attacks, the recall is lower for the higher FNR than other attacks.

### v2

@@ Abstract
WiFi fingerprint-based indoor localization schemes deliver highly accurate location data by matching the received signal strength indicator (RSSI) with an offline database using machine learning (ML) or deep learning (DL) models. However, over time, RSSI values degrade due to the malicious behavior of access points (APs), causing low positional accuracy due to RSSI value mismatch with the offline database. Existing literature lacks the detection of malicious APs in the online phase and mitigating their effects. This research addresses these limitations and proposes a long-term, reliable indoor localization scheme by incorporating malicious AP detection and their effect mitigation techniques. The proposed scheme uses a Light Gradient-Boosting Machine (LGBM) classifier to estimate locations and integrates simple yet efficient techniques to detect malicious APs based on online query data. Subsequently, a mitigation technique is incorporated that updates the offline database and online queries by imputing stable values for malicious APs using LGBM Regressors. Additionally, we introduce a noise addition mechanism in the offline database to capture the dynamic environmental effects. Extensive experimental evaluation shows that the proposed scheme attains a detection accuracy above 95% for each attack type. The mitigation strategy effectively restores the system’s performance nearly to its original state when no malicious AP is present. The noise addition module reduces localization errors by nearly 16%. Furthermore, the proposed solution is lightweight, reducing the execution time by approximately 94% compared to the existing methods.
Cite as:
F. T. Shifat, S. S. Ela and M. Jahan, “Securing WiFi Fingerprint-based Indoor Localization Systems from Malicious Access Points,” in 2025 IEEE 7th Symposium on Computers & Informatics (ISCI), Kuala Lumpur, Malaysia, 2025, pp. 380-387, doi: 10.1109/ISCI65687.2025.11167252.
Link to the final, published paper in IEEE Xplore:
https://doi.org/10.1109/ISCI65687.2025.11167252
I Introduction
With the rapid advancement of wireless communication technologies and the Internet of Things (IoT), location-based services have become indispensable in many application domains. Various localization mechanisms that operate either indoors or outdoors are employed to determine location. In outdoor environments, positions can be accurately determined using the Global Navigation Satellite System (GNSS). However, indoor positioning is challenging due to the attenuation of the Global Positioning System (GPS) signal caused by building structures and their materials [1]. Accurate indoor location tracking is critical for several purposes, such as firefighter tracking under hazardous conditions and in emergencies, location-based smart resource management, robotics, shopping mall navigation, healthcare, and virtual reality [1], [2], [3]. Consequently, indoor localization has become a major focus for both academia and industry.
@@ I Introduction
A significant security issue is the presence of malicious access points that generate false RSSI signals [8]. Over time, APs may behave maliciously due to component malfunctions [9, 10] and security attacks such as signal jamming [8], spoofing [6], distance fraud [11], [12], and physical attacks [5]. These malicious APs mislead localization systems and reduce location accuracy. Unreliable location accuracy can have serious consequences in critical situations. For example, if a person needs immediate medical attention in a hospital, inaccurate location prediction of healthcare professionals and facilities can elevate health risks. Similarly, robots navigating in complex indoor environments with unreliable localization methods may cause damage to assets and potentially lead to injuries. Existing literature deals with malicious APs during the initial database construction phase (offline phase) by selecting a subset of feasible APs [5], [6]. However, filtering out malicious APs only in the offline phase does not guarantee the long-term robustness of the indoor localization systems, as a reliable AP may become malicious over time. Current literature lacks research on detecting the long-term malicious behavior of APs [10, 13, 14, 15].
After detecting malicious APs, it is crucial to mitigate their effect to ensure the accurate functioning of the indoor localization system. One approach to achieving this goal is to reconstruct the database. However, each time a malicious AP is detected, conducting a site survey to build the database in a dynamic environment is highly expensive and labor-intensive [13]. Existing literature has explored the use of crowdsourcing techniques [15] for creating and updating the fingerprint database, which significantly reduces human labor. However, this scheme is vulnerable to various security issues [16]. Additionally, robots are utilized for autonomous database construction, which can be expensive and may not perform effectively in challenging environments [17]. Instead, it would be more beneficial to develop a mechanism that enables an indoor localization system to function accurately despite the presence of a reasonable percentage of malicious APs, eliminating the need for frequent database reconstruction. This would enhance the system’s accuracy while reducing the costs associated with repeated database creation. Although several existing research works mitigate the effect of RSSI value alteration due to environmental factors [18], [15], [13], none of them consider the malicious behavior of APs causing abrupt and dynamic changes to RSSI values.
@@ II Related works
WiFi RSSI-based indoor localization schemes provide highly accurate location data at a low additional cost. However, the location predicted by these models often becomes erroneous due to the alteration of RSSI values caused by malicious AP behaviour [5], [11], [8], changes in layout [15], removal of AP [14], position change of AP [14], [15], and environmental factors [4, 18, 14]. Existing literary works [5, 6, 19, 20] detect and filter malicious APs during the model training phase (offline phase) of indoor localization systems. Here, a subset of optimal AP is selected, and using these selected nonmalicious APs, a fingerprint database is constructed to train the machine learning model. The selection of correlated APs increases localization accuracy and reduces computational complexity. Wang et al. [5] utilized interclass dispersion to calculate AP confidence and selected APs with high confidence. Ye et al. [6] used the Pearson Correlation Coefficient (PCC), where a lower correlation value indicates greater susceptibility of an AP to being malicious. Chen et al. [19] considered the resolution capability to select a set of APs, where APs are chosen based on information gain. Panja et al. [20] used the Binary Particle Swarm Optimization (BPSO) to select a set of suitable APs. In contrast, our proposed scheme uses Spearman’s Correlation Coefficient (SRCC) to select correlated APs in the offline phase. We use normalization, correlated AP selection, and noise addition as data preprocessing, where normalization speeds up the computational time, and noise addition makes the system robust to long-term signal variation due to dynamic environmental change.
@@
Detecting malicious APs only in the offline phase is not sufficient for long-term model reliability. Several works explore long-term model robustness challenges in the online phase. For example, Yan et al. [18] proposed a crowdsourcing-based mechanism to handle RSSI signal value alteration in the online phase. This scheme utilizes a denoising autoencoder to address RSSI alteration caused by environmental dynamics. It requires extensive, high-quality crowdsourced data for training, and the model is updated frequently, which can be resource-intensive. Li et al. [10] addressed signal variation caused by dynamic environmental change and temporal effects using disagreement-based semi-supervised learning. This scheme struggles to handle sudden and significant alterations in the environment, which can impact localization accuracy. Tiku and Pasricha [4] utilized Siamese neural encoders to address RSSI signal variation without the need for model retraining. This scheme handles signal alteration caused by environmental changes, human movement, and the removal or replacement of AP. Significant changes in the environment can affect localization accuracy. Huang et al. [13] addressed location accuracy degradation caused by environmental changes. This scheme utilizes a Gaussian Process Regression to create an offline database and incorporates the Marginalized Particle Extended Gaussian process that uses crowdsourced data to recursively update the database. However, this scheme is susceptible to localization errors due to rapid fluctuations in signal strength, and its effectiveness depends on the quality of the crowdsourced data. Jiang et al. [14] addressed RSSI signal alteration caused by the change in the number of AP and environmental factors. The authors proposed a feature-adaptive extreme learning machine-based approach that can adapt to a changing number of features. This scheme faces difficulties in coping with sudden environmental changes and runs the risk of overfitting to recent signal variations. Yang et al. [15] proposed the Altered AP Identification and Fingerprint Updating (AAIFU) scheme, which detects altered APs in the online phase by considering changes in RSSI values due to positional changes of APs. This scheme uses a Gradient Boosting Decision Tree (GBDT) regression model to identify the positional changes of APs. It then updates the database by replacing the outdated RSSI values with fresh RSSI values obtained through GBDT regression. However, the effectiveness of this scheme depends on a large volume of crowdsourced data. Moreover, positional changes of an AP only create persistent changes in RSSI values. Among these works, Li et al. [10], Huang et al. [13], Jiang et al. [14], and Yang et al. [15] update the database to keep the fingerprints fresh. On the other hand, none of the schemes, except Yang et al. [15], identify the causes of RSSI signal alteration. In contrast to the existing works, our proposed scheme addresses the RSSI value alteration caused by malicious AP behavior, characterized by abrupt and random changes in RSSI values. Our proposed scheme detects malicious APs in real time by analyzing data from various location queries. Additionally, it utilizes the LGBM regression model [21] to update the existing database, which mitigates the effect of malicious APs on the performance of the indoor localization scheme. Here to note that the proposed scheme handles both static and dynamic changes in RSSI values.
@@ III System Model
As shown in Fig. 1, the proposed scheme augments the existing WiFi fingerprint-based indoor localization system [22] by including a malicious effect mitigation module and online database. The proposed system model comprises access points and reference points (RPs). The offline database comprises a set of records of form , where is the RSSI vector and is the RSSI value of th AP measured at th RP, and is the location of the th RP, , and . In the online phase, a device measures a fingerprint vector with respect to its location and sends it to the location prediction module. The location prediction module matches with the records stored in the offline database using either an ML or DL model and returns the estimated location to the mobile device. The online database is constructed by saving the results of online queries along with associated RSSI vectors and used to detect malicious APs in the online phase. The malicious effect mitigation module detects malicious APs and takes measures to diminish their effect. The proposed system model subjects to false fingerprints both in online and offline phases due to various environmental factors [4, 14], temporal effects [6], and malicious behavior of APs [5, 6]. We assume that the databases, the location prediction module, and the malicious effect mitigation module reside in a trusted server.
@@ IV-A1 Database Construction
At first, a reliable database is created using the site survey [23]. This mechanism involves a person with multiple reliable mobile devices visiting every RP. It collects several labeled samples in the form , where is an RSSI fingerprint vector computed at th RP and is the location of the th RP. The raw database created by the site survey is a collection of records .
@@ IV-A2 Correlated AP Selection
The proposed scheme selects a set of reliable APs from the available ones in the environment using Spearman’s Rank Correlation Coefficient (SRCC) [24] method with a threshold value set to 0.1 empirically, which is optimal for the considered UJI’s dataset [25]. Suppose the number of selected correlated APs is . The proposed scheme constructs an offline database from the raw database constructed via site survey, where contains RSSI entries only for the selected APs.
@@ IV-C1 Check Variance
It checks the variance of an AP for every RP in the online database. If the maximum variance value of an AP equals or exceeds the variance threshold, the AP is marked as malicious. The variance threshold is set to 0.05, as nearly 99% of the variance for each AP in the offline database is within 0.05 for the UJI dataset [25]. Let,
@@ IV-D2 Update Online Queries and Offline Database
The LGBMRegressors predicts RSSI values when input data is of form and use these data to replace the values of malicious APs , in the offline record. Similarly, when the input data is of form , the predicted RSSI values replace , in the online query . The localization model is retrained with the updated offline database. The updated query is used by the classification model to predict the location. The process of predicting RSSI values is inspired by Montoliu et al. [27]. The malicious AP detection module is executed after a fixed interval determined by the online database size, number of users, time to serve a location query and the number of requests/minute. If a malicious AP is detected, then the malicious effect mitigation module is executed.
@@ V Experimental Evaluation
We implemented the proposed scheme using Python and evaluated its performance using the UJI’s dataset [25]. The UJI’s dataset [25] is the longest public dataset with 103584 WiFi fingerprints. We used only the data of floor 3 for brevity, as also done in [4]. This data set contains 4320 data samples for training purposes for floor 3 [28]. We incorporated noise-added samples in this set and extended its size to 4800. Initially, there were 620 APs in the environment. However, we took only 40 APs after correlated AP selection. We trained our localization model which is an LGBMClassifier [21], with the offline dataset. We used seven performance metrics shown in Table III to evaluate the performance. The parameters used in defining performance metrics are shown in Table II.
@@ V-A Impact of Noise Addition
Figure 3 shows mean localization errors over the data of 25 months from UJI’s dataset [25]. The curve labeled SRCC (0.1) shows the effect of correlated AP selection only, the curve labeled SRCC (0.1) + Normalized shows the impact of AP selection and normalization, and the curve labeled SRCC (0.1) + Normalized + Noise (0.5) shows the effect of AP selection, normalization, and noise addition. Figure 3 shows a notable reduction in localization error for the curve SRCC (0.1) + Normalized + Noise (0.5) with respect to SRCC (0.1) + Normalized after the 10th month because the model becomes robust to RSSI fluctuations in the long term. The MLE is reduced by nearly 16% in the 16th month, which is the maximum change of the red curve with respect to the green one. Normalizing data helps models converge faster, and adding noise makes the model more robust against long-term noise. Correlated AP selection, noise addition, and data normalization bring the MLE of the proposed scheme below 2 meters, outperforming existing machine learning schemes [4].
@@ V-B1 Performance Analysis of the Proposed Scheme
False Positive Rate (FPR): Figure 4(a) shows that the FPR consistently remains below 4%. The legitimate APs sometimes exhibit signal variation due to temporal effects and are marked as malicious by the detection system.
@@
False Negative Rate (FNR): Figure 4(b) shows that, except for the ARCO attack, all other attacks show a negligible FNR. As marking an AP as malicious in our experiment was a random process, in very few cases, the selected malicious AP was assigned the same original RSSI value, leading to a false negative outcome. In ARCO attacks, FNR stays nearly 10%. In this attack, we add or subtract a fixed value and also add an offset in the noise addition module, leading to very little change in the mean that can not be distinguished with the mean difference checker, resulting in an increased FNR.
@@
Recall: The recall is consistently 100% for RV and ARRO attacks, almost 98% for CV attacks, and nearly 92% for ARCO attacks, as shown in Fig. 4(c). In ARCO attacks, the recall is lower for the higher FNR than for other attacks.